OpenStudy (anonymous):

Which of the following has a solution of 4? 5 times the reciprocal of four-fifths nine divided by four-ninths six-sevenths of four and two-thirds three-fifths times three-twentieths

OpenStudy (anonymous):

5 times reciprocal of 4/5 : 5 * 5/4 = 20/4 = 5 not this one 9/(4/9) = 9 * 9/4 = 81/4 = 20 1/4 not this one 6/7 of 4 2/3 --- " of " means multiply 6/7 * 14/3 = 84/21 = 4 BINGO....found it. Let me go ahead and check the last one.. 3/5 * 3/12 = 9/60 = 3/20 It is the third one :)